5) Please Help

5. Unpolarized light with intensity lo is incident on two polarizing filters. The axis of the first filter makes an angle of 40.0° with the vertical, and the axis of the second filter is horizontal. What is the intensity of the light after it has passed through the second filter?

An unpolorized light with intensity i0 is incident on two polorizing filter. The axis of the first filter makes an angle of 40 with the verticla and the axis of the secons filter is horizontal. What is the intensity of the light after it has passed through the second filter?
Answer #1

here,

initial intensity of light is i0

theta = 40 degree

when light passes through first polarizer , i1 = 0.5 * i0

when the light is passed throught the seccond polarizer , i2 = i1 * ( cos(phi))^2

i2 = 0.5 * i0 * ( cos(90 - theta))^2

i2 = 0.5 * i0 * ( cos(90 - 40))^2

i2 = 0.21 * i0

the intensity of light passed through the seccond polarizer is 0.21 * i0